OpenStudy (anonymous):

I'm still kind of having problems with exponents. I really don't understand when there are fractions involved! (1/4)^(2/3) or 8b^(1/3) Please show work and thank you!

OpenStudy (anonymous):

Also when there is negatives involved! Here is an example from my work: (125/64)^(-1/3)

OpenStudy (ybarrap):

Let's take the first one: $$ \Large{ 1^\frac{2}{3} = (1^{1/3})^2=(\sqrt[3]{1})^2\\ 4^\frac{2}{3} = (4^{1/3})^2=(\sqrt[3]{4})^2\\ \left( \cfrac{1}{4}\right )^{2/3}=\cfrac{1^\frac{2}{3}}{4^\frac{2}{3}}=\cfrac{(\sqrt[3]{1})^2}{(\sqrt[3]{4})^2}=\left (\sqrt[3]{\cfrac{1}{4}}\right )^{2} } $$ Does this make sense?

OpenStudy (anonymous):

Yeah(: Thanks, but my teacher doesn't like it when we leave is in that format. He likes when they're fractions (without exponents) and can you do the last example I just added?

OpenStudy (anonymous):

A negative exponent is just 1/whatever. Like \(x^{-2}\) is just \(\frac{1}{x^2}\)

OpenStudy (anonymous):

\[\large \left(\frac{125}{64}\right)^{-\frac{1}{3}}\] the minus sign in the exponent means flip it the 3 in the denominator means take the cubed root

OpenStudy (anonymous):

so 5/4?

OpenStudy (anonymous):

Do you mean taking the exponent of a fraction? Just flip the fraction over.

OpenStudy (anonymous):

The negative exponent, I mean.

OpenStudy (ybarrap):

$$ \Large{ (8b)^{1/3}=\sqrt[3]{8b} } $$

OpenStudy (anonymous):

close but you forgot to flip it

OpenStudy (anonymous):

@ybarrap Could the answer be 4 times the cubed root of b?

OpenStudy (ybarrap):

$$ \Large{ (8b)^{1/3}=\sqrt[3]{8b}=2\sqrt[3]{b} } $$

OpenStudy (anonymous):

work through the worksheet i sent, i bet if you do it step by step you will get it. look at the examples first

OpenStudy (anonymous):

OH you flip the inside number not the exponent?

OpenStudy (anonymous):

right!!

OpenStudy (anonymous):

no one thinks for example that \(-\frac{2}{3}=\frac{3}{2}\) if the exponent is negative, the notation means take the reciprocal (flip it)

OpenStudy (anonymous):

Okay(: Um what if it's an integer and a fractional exponent? like: (25+144)^1/2

OpenStudy (anonymous):

\[(\frac{a}{b})^{-1} = \frac{1}{\frac{a}{b}} = \frac{b}{a}\]

OpenStudy (anonymous):

then don't flip it leave it alone

OpenStudy (anonymous):

Wait, it's the negative reciprocal of the exponent or inside number? @satellite73

OpenStudy (anonymous):

just the reciprocal of the number inside don't make it negative ever

OpenStudy (anonymous):

Okay thanks(: @satellite73

OpenStudy (anonymous):

\[\large 27^{-\frac{1}{3}}=\frac{1}{\sqrt[3]{27}}=\frac{1}{3}\] for example

OpenStudy (anonymous):

Wow everyone is so nice(: Thanks! So when the inside number and exponents are fractions you just distribute the exponent and solve from then on?

OpenStudy (ranga):

\[\Large (\frac{ a }{ b })^{-\frac{ c }{ d }} = (\frac{ b }{ a })^{+\frac{ c }{ d }}\]

OpenStudy (anonymous):

Okay, wow this is really hard. 32^(-3/5)?

OpenStudy (ybarrap):

$$ \Large{ \left (\cfrac{125}{64}\right )^{-1/3}\\ =\left (\cfrac{125^{-1}}{64^{-1}}\right )^{1/3}\\ =\left (\cfrac{64}{125}\right )^{1/3}\\ =\sqrt[3]{\cfrac{64}{125}}\\ =\cfrac{4}{?} } $$

OpenStudy (anonymous):

? = 5 right? @ybarrap

OpenStudy (ybarrap):

yes

OpenStudy (ybarrap):

$$ \Large{ 32^{-3/5}\\ (32^{-1})^{3/5}\\ \left (\cfrac{1}{32}\right )^{3/5}\\ =\cfrac{1}{(32)^{3/5}}\\ =\cfrac{1}{(\sqrt[5]{3})^3} } $$

OpenStudy (anonymous):

There's no way to simplify that any more?

OpenStudy (ybarrap):

If \(\large(\sqrt[5]{3})^3=\sqrt[5]{3}\times\sqrt[5]{3}\times\sqrt[5]{3}\) can be simplified, then yes. But I don't see any way.

OpenStudy (anonymous):

Okay! wow thanks everyone! (: I actually get it now (surprise)

OpenStudy (ybarrap):

yw